# https://kaist-viclab.github.io/SLiM_site/

### 11 Mar 2026 v1

@@ 4.3 Performance Comparison
We evaluate our SLiM against state-of-the-art self-supervised approaches, including CL, MAE, and other pretext tasks [14, 37, 16, 36, 35]. To ensure a fair comparison and isolate the efficacy of the learned representations, we report results based on the single joint modality across all baselines, excluding gains from multi-stream fusion (e.g., bone or motion modalities).
@@
GL-Transformer [14] ECCV’22 76.3 83.8 66.0 68.7
MacDiff [37] ECCV’24 86.4 91.0 79.4 80.2
IGM [16] ECCV’24 86.2 91.2 80.0 81.4
USDRL [36] AAAI’25 85.2 91.7 76.6 78.1 54.4
HSP [35] CVPR’25 80.7 88.0 71.0 73.2 48.9
@@
CPM [39] ECCV’22 78.7 84.9 68.7 69.6 48.3
CMD [23] ECCV’22 79.8 86.9 70.3 71.5 43.0
AimCLR [11] AAAI’22 74.3 79.7 63.4 63.4
RVTCLR [44] ICCV’23 74.7 79.1
PTSL [43] AAAI’23 77.3 81.8 66.2 67.7 49.3
HYSP [10] ICLR’23 78.2 82.6 61.8 64.6
HaLP [25] CVPR’23 79.7 86.8 71.1 72.2 43.5
ActCLR [17] CVPR’23 80.9 86.7 69.0 70.5
@@
SkeletonMAE [38] ICCV’23 74.8 77.7 72.5 73.5 36.1
MAMP [22] ICCV’23 84.9 89.1 78.6 79.1 53.8
S-JEPA [1] ECCV’24 85.3 89.8 79.6 79.9 53.5
GFP [32] ICCV’25 85.9 92.0 79.1 80.3 56.2
@@
GL-Transformer [14] 68.6 74.9
CPM [39] 56.7 73.0 57.5 77.1
CMD [23] 50.6 75.4 53.0 80.2
HYSP [10] 76.2 80.4
HaLP [25] 46.6 72.6 48.7 77.1
HiCo [7] 54.4 73.0 54.8 78.3
UmURL [31] 58.1 58.3
USDRL [36] 57.3 80.2 60.7 84.0
SkeletonMAE [38] 54.4 80.6 54.6 83.5
MAMP [22] 66.0 88.0 68.7 91.5
S-JEPA [1] 67.5 88.4 69.1 91.4
GFP [32] 71.8 88.7 72.9 92.1
@@
LongT-GAN [41] 39.1 48.1
P&C [28] 50.7 76.3
ISC [29] 62.5 82.6
HaLP [25] 65.8 83.6
HiCo [7] 68.3 84.8
SkeAttnCLR [13] 69.4 76.8
UmURL [31] 71.3 88.3
MAMP [22] 62.0 70.0
GFP [32] 70.9 87.1
@@
Action recognition (linear evaluation). Table 4.3 reports the Top-1 accuracy under the standard linear evaluation [22, 1, 32]. In this setting, we freeze the parameters of the pre-trained encoder and train only a single-layer linear classifier for evaluation. While previous methods often showed inconsistent performance across different benchmarks, SLiM establishes a new state-of-the-art across multiple evaluated datasets: NTU-60, NTU-120, and PKU-MMD II. As shown in Table 4.3, our framework consistently outperforms all existing representation learning paradigms. Remarkably, as highlighted in Table 1, our performance gains are achieved with significantly reduced computational overhead, requiring only 3.59 GFLOPs for inference—a $7.89\times$ reduction in computational cost compared to traditional MAE methods. Specifically, on the large-scale NTU-120 benchmark, SLiM achieves 81.2% on X-Sub and 83.6% on X-Set, surpassing the previous best method by 1.2%-point and 2.2%-point, respectively. On NTU-60, our method demonstrates superior robustness with 87.9% on X-Sub and 93.2% on X-View, representing a significant lead over existing frameworks. Furthermore, SLiM sets a new record on the PKU-MMD II dataset with 59.7% on X-Sub. Across the five primary evaluation protocols, SLiM provides an average accuracy improvement of approximately 1.9%-point over the closest competitors.
Action recognition (semi-supervised learning). To evaluate the data efficiency and label robustness of the pre-trained representations, we conduct semi-supervised learning experiments on the NTU-60 dataset. Table 4 shows the performance comparison for action recognition under 1% and 10% label settings. For the experimental setup, we follow the standard evaluation protocol [22, 1, 32] by fine-tuning the pre-trained encoder using only these limited labeled subsets and measure the performance according to NTU-60. SLiM significantly outperforms MAE methods under the extremely label-sparse 1% setting, while the performance gap narrows at 10%. This phenomenon stems from differences in sequence length, token density, and network capacity. Previous MAE models process $T=120$ frames with a fine-grained temporal patch size of $P_{T}=4$, resulting in extended token sequences that require extensive data to properly optimize. In contrast, SLiM utilizes $T=64$ frames with a coarser patch size of $P_{T}=8$, operating on merely $\sim$25% of the tokens. Under extreme label sparsity (1%), this compact representation, combined with our strong geometric priors, effectively prevents overfitting and enables rapid convergence. When supervision increases to 10%, the additional data provides sufficient gradient signals for heavier models to leverage the higher capacity of their dense token spaces, marginally closing the performance gap. Nevertheless, SLiM still achieves strictly comparable or superior results while requiring a drastically reduced computational footprint.
@@ 4.4 Ablation Studies
Efficacy of dual-role STM. Table 6 demonstrates the impact of our masking strategies. When applying the independent joint masking common in prior works [22, 1, 32], the model easily cheats the pretext task via spatial interpolation from adjacent coordinates. Consequently, utilizing this limited strategy yields suboptimal performance: 84.2% on X-Sub for the global view ($\mathcal{L}_{\text{MFM}}$ target) and 85.7% for the local view ($\mathcal{L}_{\text{GLCL}}$ perturbation). In contrast, our STM prevents such naive shortcut learning by masking entire anatomical parts across the temporal axis. STM serves a dual role: providing robust targets for $\mathcal{L}_{\text{MFM}}$ and acting as a strong structural perturbation for $\mathcal{L}_{\text{GLCL}}$. By driving the encoder to deeply understand motion dynamics, integrating STM across both objectives achieves a peak performance of 87.7%, demonstrating substantial improvements (+3.5%-point and +2.0%-point, respectively) over conventional masking.
Impact of SAA. Table 7 isolates the contributions of geometric augmentations. By substituting standard augmentations from prior works [17, 11] with our SAA—encompassing skeleton-aware rotation, mirroring, and bone-aware scaling—we strictly preserve the 3D geometric validity of contrastive pairs. This anatomical consistency directly translates to enhanced downstream accuracy. Consequently, compared to their conventional counterparts, our rotation, mirroring, and scaling yield substantial improvements of +2.9%-point, +1.8%-point, and +1.3%-point on X-Sub, respectively.
@@ Acknowledgements
Transfer learning evaluation on PKU-MMD II [18] (Table 8)
In this Appendix, we provide the detailed algorithmic formulations for the data augmentation and masking strategies employed in our SLiM framework. Our pipeline consists of a dual-role Semantic Tube Masking (STM) and a suite of improved Skeletal-Aware Augmentations (SAA) specifically designed to ensure anatomical consistency across diverse temporal granularities. Furthermore, we detail our hierarchical temporal sampling strategy, which anchors short-term local clips within long-term global segments to maintain semantic consistency. This approach, combined with our decoder-free architecture, allows SLiM to learn robust and transferable skeletal representations with high computational efficiency.
@@ Appendix 0.A Implementation Details
Detailed architecture. Our encoder $g$ follows a ViT [8] variant optimized for skeleton data, consisting of 8 Transformer layers with a hidden dimension of $D=256$ and 8 attention heads. To enhance the stability of attention maps and prevent feature collapse, we introduce 4 register tokens ($N_{\text{reg}}=4$) [5], which allows the model to isolate outlier features. Both student network $f_{\bm{\theta}}$ and teacher network $f_{\bm{\phi}}$ utilize identical 3-layer MLP projection heads ($h_{\texttt{CL}}$ and $h_{\texttt{MAE}}$) with a hidden dimension of 2,048 and a bottleneck dimension of 256. These heads project features onto a high-dimensional prototype space of $K=65,536$ dimensions to facilitate fine-grained feature discrimination. Consequently, the probability distributions $\mathbf{p}$ in Eq. 4 and Eq. 6 are defined over this $K$-dimensional categorical space, forcing the model to distinguish between tens of thousands of refined motion patterns.
Loss functions. The total loss is optimized by balancing the masked feature modeling loss $\mathcal{L}_{\text{MFM}}$ and the global-local contrastive loss $\mathcal{L}_{\text{GLCL}}$ with an equal weight of 1:1, setting the hyperparameter $\lambda=1$ in Eq. 7. Additionally, we incorporate the KoLeo loss [24] with a weight of $0.1$ to encourage the local features to be spread uniformly across the unit sphere, further enhancing the discriminative power of the learned representations. For prototype centering, we adopt the Sinkhorn-Knopp algorithm [42, 24] to ensure a uniform distribution of assignments across the prototype space.
@@ Appendix 0.C Semantic Tube Masking (STM)
Concept and motivation. As shown in Fig. 3, independent joint masking [38, 22, 1, 32] often permits the model to bypass robust representation learning by simply interpolating missing coordinates from immediate spatial neighbors. To prevent this trivial shortcut reconstruction, Semantic Tube Masking (STM) explicitly masks integral anatomical groups (e.g., arms, legs, or torso) continuously across the temporal dimension, thereby forming continuous skeletal-temporal tubes.
Anatomical joint grouping. To implement this on standard 25-joint skeleton sequences (e.g., the Kinect v2 topology utilized in the NTU-60 [26], NTU-120 [19], and PKU-MMD II [18] datasets), we define the set of semantic joint groups $\mathcal{G}$ by dividing the human body into five distinct anatomical regions: Torso $G_{1}=\{0,1,20,2,3\}$, Left Arm $G_{2}=\{4,5,6,7,21,22\}$, Right Arm $G_{3}=\{8,9,10,11,23,24\}$, Left Leg $G_{4}=\{12,13,14,15\}$, and Right Leg $G_{5}=\{16,17,18,19\}$. To ensure that masking does not degrade into independent joint occlusion within these regions, we randomly select a anatomically contiguous (i.e., skeletally connected) subset of joints $S\subseteq G$ to form the spatial mask. For example, rather than dropping disconnected joints from the Left Arm $G_{2}$, we select a contiguous sequence (e.g., the elbow, wrist, and hand joints $\{5,6,7\}$), guaranteeing that a connected spatial block is occluded and preventing trivial interpolation.
@@ Appendix 0.D Skeleton-Aware Augmentations (SAA)
As illustrated in Fig. 3, standard geometric transformations [17, 11, 45] frequently overlook the physical constraints of articulated human bodies, resulting in distorted limb proportions or physically implausible postures. Our SAA framework addresses this limitation by constructing diverse yet structurally valid contrastive pairs. To ensure sufficient view diversity while maintaining anatomical integrity, each augmentation within the SAA suite is applied independently with a probability of $p=0.5$.
@@ 0.D.1 Skeleton-Aware Rotation (Algorithm 2)
Traditional uniform rotation strategies [17, 4, 11] are typically restricted to narrow angle ranges to avoid unrealistic poses. Our approach explicitly decouples the rotation logic. Specifically, we sample the vertical rotation angle $\beta$ from a uniform distribution $\mathcal{U}(-\theta_{\text{vert}},\theta_{\text{vert}})$ with $\theta_{\text{vert}}=180^{\circ}$, effectively applying a full $360^{\circ}$ random rotation exclusively around the vertical $Y$-axis to preserve a realistic upright posture. Conversely, the tilt angles $\alpha$ and $\gamma$ along the non-gravity axes ($X$ and $Z$, respectively) are sampled from $\mathcal{U}(-\theta_{\text{tilt}},\theta_{\text{tilt}})$. We bound these perturbations to small angles by setting $\theta_{\text{tilt}}=30^{\circ}$ to prevent physically impossible tilting.
@@ 0.D.2 Skeleton-Aware Mirroring (Algorithm 3)
Conventional index swapping [17, 11, 6] fails to execute a mathematically rigorous reflection, leading to unnatural motion artifacts. Our geometric mirroring strategy simultaneously negates the spatial coordinate values along the lateral $X$-axis and reassigns the corresponding left-right joint indices via a predefined bilateral mapping $\mathcal{I}_{\text{swap}}$. Specifically, for the standard 25-joint Kinect v2 topology utilized in the NTU-60 [26], NTU-120 [19], and PKU-MMD II [18] datasets, $\mathcal{I}_{\text{swap}}$ explicitly swaps the symmetric anatomical pairs across the body’s vertical midline: the arms $\{(5,9),(6,10),(7,11),(8,12)\}$, the legs $\{(13,17),(14,18),(15,19),(16,20)\}$, and the hands $\{(22,24),(23,25)\}$, while central spine and head joints remain at their original indices. This dual operation—spatial negation coupled with rigorous index swapping—guarantees that the left-right semantics and the correct “Front” orientation are anatomically preserved.

### 12 Mar 2026 v2

@@ 4.3 Performance Comparison
We evaluate our SLiM against state-of-the-art self-supervised approaches, including CL, MAE, and other pretext tasks [GL-Transformer, MacDiff, IGM, USDRL, HSP]. To ensure a fair comparison and isolate the efficacy of the learned representations, we report results based on the single joint modality across all baselines, excluding gains from multi-stream fusion (e.g., bone or motion modalities).
@@
GL-Transformer [GL-Transformer] ECCV’22 76.3 83.8 66.0 68.7
MacDiff [MacDiff] ECCV’24 86.4 91.0 79.4 80.2
IGM [IGM] ECCV’24 86.2 91.2 80.0 81.4
USDRL [USDRL] AAAI’25 85.2 91.7 76.6 78.1 54.4
HSP [HSP] CVPR’25 80.7 88.0 71.0 73.2 48.9
@@
CPM [CPM] ECCV’22 78.7 84.9 68.7 69.6 48.3
CMD [CMD] ECCV’22 79.8 86.9 70.3 71.5 43.0
AimCLR [AimCLR] AAAI’22 74.3 79.7 63.4 63.4
RVTCLR [RVTCLR] ICCV’23 74.7 79.1
PTSL [PTSL] AAAI’23 77.3 81.8 66.2 67.7 49.3
HYSP [HYSP] ICLR’23 78.2 82.6 61.8 64.6
HaLP [HaLP] CVPR’23 79.7 86.8 71.1 72.2 43.5
ActCLR [ActCLR] CVPR’23 80.9 86.7 69.0 70.5
@@
SkeletonMAE [SkeletonMAE] ICCV’23 74.8 77.7 72.5 73.5 36.1
MAMP [MAMP] ICCV’23 84.9 89.1 78.6 79.1 53.8
S-JEPA [SJEPA] ECCV’24 85.3 89.8 79.6 79.9 53.5
GFP [GFP] ICCV’25 85.9 92.0 79.1 80.3 56.2
@@
GL-Transformer [GL-Transformer] 68.6 74.9
CPM [CPM] 56.7 73.0 57.5 77.1
CMD [CMD] 50.6 75.4 53.0 80.2
HYSP [HYSP] 76.2 80.4
HaLP [HaLP] 46.6 72.6 48.7 77.1
HiCo [HiCo] 54.4 73.0 54.8 78.3
UmURL [UmURL] 58.1 58.3
USDRL [USDRL] 57.3 80.2 60.7 84.0
SkeletonMAE [SkeletonMAE] 54.4 80.6 54.6 83.5
MAMP [MAMP] 66.0 88.0 68.7 91.5
S-JEPA [SJEPA] 67.5 88.4 69.1 91.4
GFP [GFP] 71.8 88.7 72.9 92.1
@@
LongT-GAN [LongTGAN] 39.1 48.1
P&C [PandC] 50.7 76.3
ISC [ISC] 62.5 82.6
HaLP [HaLP] 65.8 83.6
HiCo [HiCo] 68.3 84.8
SkeAttnCLR [SkeAttnCLR] 69.4 76.8
UmURL [UmURL] 71.3 88.3
MAMP [MAMP] 62.0 70.0
GFP [GFP] 70.9 87.1
@@
Action recognition (linear evaluation). Table 4.3 reports the Top-1 accuracy under the standard linear evaluation [MAMP, SJEPA, GFP]. In this setting, we freeze the parameters of the pre-trained encoder and train only a single-layer linear classifier for evaluation. While previous methods often showed inconsistent performance across different benchmarks, SLiM establishes a new state-of-the-art across multiple evaluated datasets: NTU-60, NTU-120, and PKU-MMD II. As shown in Table 4.3, our framework consistently outperforms all existing representation learning paradigms. Remarkably, as highlighted in Table 1, our performance gains are achieved with significantly reduced computational overhead, requiring only 3.59 GFLOPs for inference—a $7.89\times$ reduction in computational cost compared to traditional MAE methods. Specifically, on the large-scale NTU-120 benchmark, SLiM achieves 81.2% on X-Sub and 83.6% on X-Set, surpassing the previous best method by 1.2%-point and 2.2%-point, respectively. On NTU-60, our method demonstrates superior robustness with 87.9% on X-Sub and 93.2% on X-View, representing a significant lead over existing frameworks. Furthermore, SLiM sets a new record on the PKU-MMD II dataset with 59.7% on X-Sub. Across the five primary evaluation protocols, SLiM provides an average accuracy improvement of approximately 1.9%-point over the closest competitors.
Action recognition (semi-supervised learning). To evaluate the data efficiency and label robustness of the pre-trained representations, we conduct semi-supervised learning experiments on the NTU-60 dataset. Table 4 shows the performance comparison for action recognition under 1% and 10% label settings. For the experimental setup, we follow the standard evaluation protocol [MAMP, SJEPA, GFP] by fine-tuning the pre-trained encoder using only these limited labeled subsets and measure the performance according to NTU-60. SLiM significantly outperforms MAE methods under the extremely label-sparse 1% setting, while the performance gap narrows at 10%. This phenomenon stems from differences in sequence length, token density, and network capacity. Previous MAE models process $T=120$ frames with a fine-grained temporal patch size of $P_{T}=4$, resulting in extended token sequences that require extensive data to properly optimize. In contrast, SLiM utilizes $T=64$ frames with a coarser patch size of $P_{T}=8$, operating on merely $\sim$25% of the tokens. Under extreme label sparsity (1%), this compact representation, combined with our strong geometric priors, effectively prevents overfitting and enables rapid convergence. When supervision increases to 10%, the additional data provides sufficient gradient signals for heavier models to leverage the higher capacity of their dense token spaces, marginally closing the performance gap. Nevertheless, SLiM still achieves strictly comparable or superior results while requiring a drastically reduced computational footprint.
@@ 4.4 Ablation Studies
Efficacy of dual-role STM. Table 6 demonstrates the impact of our masking strategies. When applying the independent joint masking common in prior works [MAMP, SJEPA, GFP], the model easily cheats the pretext task via spatial interpolation from adjacent coordinates. Consequently, utilizing this limited strategy yields suboptimal performance: 84.2% on X-Sub for the global view ($\mathcal{L}_{\text{MFM}}$ target) and 85.7% for the local view ($\mathcal{L}_{\text{GLCL}}$ perturbation). In contrast, our STM prevents such naive shortcut learning by masking entire anatomical parts across the temporal axis. STM serves a dual role: providing robust targets for $\mathcal{L}_{\text{MFM}}$ and acting as a strong structural perturbation for $\mathcal{L}_{\text{GLCL}}$. By driving the encoder to deeply understand motion dynamics, integrating STM across both objectives achieves a peak performance of 87.7%, demonstrating substantial improvements (+3.5%-point and +2.0%-point, respectively) over conventional masking.
Impact of SAA. Table 7 isolates the contributions of geometric augmentations. By substituting standard augmentations from prior works [ActCLR, AimCLR] with our SAA—encompassing skeleton-aware rotation, mirroring, and bone-aware scaling—we strictly preserve the 3D geometric validity of contrastive pairs. This anatomical consistency directly translates to enhanced downstream accuracy. Consequently, compared to their conventional counterparts, our rotation, mirroring, and scaling yield substantial improvements of +2.9%-point, +1.8%-point, and +1.3%-point on X-Sub, respectively.
@@ Acknowledgements
Transfer learning evaluation on PKU-MMD II [PKU] (Table 8)
This Appendix provides comprehensive technical and algorithmic details for the SLiM framework. We begin in Sec. 0.A by elaborating on the implementation details, including the specific network architecture, loss function formulations, and hyper-parameter configurations used throughout our experiments. In Sec. 0.B, we describe the hierarchical temporal sampling strategy, which explains the generation of global and local views to maintain semantic consistency. The formal algorithmic procedures for our core modules are then detailed: Sec. 0.C presents the Semantic Tube Masking (STM), while Sec. 0.D provides the suite of Skeleton-Aware Augmentations (SAA), specifically covering rotation, mirroring, and scaling algorithms. Finally, Sec. 0.E concludes with additional results and discussions, including transfer learning evaluations.
@@ Appendix 0.A Implementation Details
Detailed architecture. Our encoder $g$ follows a ViT [ViT] variant optimized for skeleton data, consisting of 8 Transformer layers with a hidden dimension of $D=256$ and 8 attention heads. To enhance the stability of attention maps and prevent feature collapse, we introduce 4 register tokens ($N_{\text{reg}}=4$) [Register], which allows the model to isolate outlier features. Both student network $f_{\bm{\theta}}$ and teacher network $f_{\bm{\phi}}$ utilize identical 3-layer MLP projection heads ($h_{\texttt{CL}}$ and $h_{\texttt{MAE}}$) with a hidden dimension of 2,048 and a bottleneck dimension of 256. These heads project features onto a high-dimensional prototype space of $K=65,536$ dimensions to facilitate fine-grained feature discrimination. Consequently, the probability distributions $\mathbf{p}$ in Eq. 4 and Eq. 6 are defined over this $K$-dimensional categorical space, forcing the model to distinguish between tens of thousands of refined motion patterns.
Loss functions. The total loss is optimized by balancing the masked feature modeling loss $\mathcal{L}_{\text{MFM}}$ and the global-local contrastive loss $\mathcal{L}_{\text{GLCL}}$ with an equal weight of 1:1, setting the hyperparameter $\lambda=1$ in Eq. 7. Additionally, we incorporate the KoLeo loss [DINOv2] with a weight of $0.1$ to encourage the local features to be spread uniformly across the unit sphere, further enhancing the discriminative power of the learned representations. For prototype centering, we adopt the Sinkhorn-Knopp algorithm [iBOT, DINOv2] to ensure a uniform distribution of assignments across the prototype space.
@@ Appendix 0.C Semantic Tube Masking (STM)
Concept and motivation. As shown in Fig. 3, independent joint masking [SkeletonMAE, MAMP, SJEPA, GFP] often permits the model to bypass robust representation learning by simply interpolating missing coordinates from immediate spatial neighbors. To prevent this trivial shortcut reconstruction, Semantic Tube Masking (STM) explicitly masks integral anatomical groups (e.g., arms, legs, or torso) continuously across the temporal dimension, thereby forming continuous skeletal-temporal tubes.
Anatomical joint grouping. To implement this on standard 25-joint skeleton sequences (e.g., the Kinect v2 topology utilized in the NTU-60 [NTU60], NTU-120 [NTU120], and PKU-MMD II [PKU] datasets), we define the set of semantic joint groups $\mathcal{G}$ by dividing the human body into five distinct anatomical regions: Torso $G_{1}=\{0,1,20,2,3\}$, Left Arm $G_{2}=\{4,5,6,7,21,22\}$, Right Arm $G_{3}=\{8,9,10,11,23,24\}$, Left Leg $G_{4}=\{12,13,14,15\}$, and Right Leg $G_{5}=\{16,17,18,19\}$. To ensure that masking does not degrade into independent joint occlusion within these regions, we randomly select a anatomically contiguous (i.e., skeletally connected) subset of joints $S\subseteq G$ to form the spatial mask. For example, rather than dropping disconnected joints from the Left Arm $G_{2}$, we select a contiguous sequence (e.g., the elbow, wrist, and hand joints $\{5,6,7\}$), guaranteeing that a connected spatial block is occluded and preventing trivial interpolation.
@@ Appendix 0.D Skeleton-Aware Augmentations (SAA)
As illustrated in Fig. 3, standard geometric transformations [ActCLR, AimCLR, CrosSCLR] frequently overlook the physical constraints of articulated human bodies, resulting in distorted limb proportions or physically implausible postures. Our SAA framework addresses this limitation by constructing diverse yet structurally valid contrastive pairs. To ensure sufficient view diversity while maintaining anatomical integrity, each augmentation within the SAA suite is applied independently with a probability of $p=0.5$.
@@ 0.D.1 Skeleton-Aware Rotation (Algorithm 2)
Traditional uniform rotation strategies [ActCLR, CTRGCN, AimCLR] are typically restricted to narrow angle ranges to avoid unrealistic poses. Our approach explicitly decouples the rotation logic. Specifically, we sample the vertical rotation angle $\beta$ from a uniform distribution $\mathcal{U}(-\theta_{\text{vert}},\theta_{\text{vert}})$ with $\theta_{\text{vert}}=180^{\circ}$, effectively applying a full $360^{\circ}$ random rotation exclusively around the vertical $Y$-axis to preserve a realistic upright posture. Conversely, the tilt angles $\alpha$ and $\gamma$ along the non-gravity axes ($X$ and $Z$, respectively) are sampled from $\mathcal{U}(-\theta_{\text{tilt}},\theta_{\text{tilt}})$. We bound these perturbations to small angles by setting $\theta_{\text{tilt}}=30^{\circ}$ to prevent physically impossible tilting.
@@ 0.D.2 Skeleton-Aware Mirroring (Algorithm 3)
Conventional index swapping [ActCLR, AimCLR, Skateformer] fails to execute a mathematically rigorous reflection, leading to unnatural motion artifacts. Our geometric mirroring strategy simultaneously negates the spatial coordinate values along the lateral $X$-axis and reassigns the corresponding left-right joint indices via a predefined bilateral mapping $\mathcal{I}_{\text{swap}}$. Specifically, for the standard 25-joint Kinect v2 topology utilized in the NTU-60 [NTU60], NTU-120 [NTU120], and PKU-MMD II [PKU] datasets, $\mathcal{I}_{\text{swap}}$ explicitly swaps the symmetric anatomical pairs across the body’s vertical midline: the arms $\{(5,9),(6,10),(7,11),(8,12)\}$, the legs $\{(13,17),(14,18),(15,19),(16,20)\}$, and the hands $\{(22,24),(23,25)\}$, while central spine and head joints remain at their original indices. This dual operation—spatial negation coupled with rigorous index swapping—guarantees that the left-right semantics and the correct “Front” orientation are anatomically preserved.